# Beyond Accuracy: Evaluating the Reasoning Behavior of Large Language Models - A Survey

### v1

@@ Abstract
Large language models (LLMs) have recently shown impressive performance on tasks involving reasoning, leading to a lively debate on whether these models possess reasoning capabilities similar to humans. However, despite these successes, the depth of LLMs’ reasoning abilities remains uncertain. This uncertainty partly stems from the predominant focus on task performance, measured through shallow accuracy metrics, rather than a thorough investigation of the models’ reasoning behavior. This paper seeks to address this gap by providing a comprehensive review of studies that go beyond task accuracy, offering deeper insights into the models’ reasoning processes. Furthermore, we survey prevalent methodologies to evaluate the reasoning behavior of LLMs, emphasizing current trends and efforts towards more nuanced reasoning analyses. Our review suggests that LLMs tend to rely on surface-level patterns and correlations in their training data, rather than on genuine reasoning abilities. Additionally, we identify the need for further research that delineates the key differences between human and LLM-based reasoning. Through this survey, we aim to shed light on the complex reasoning processes within LLMs.
@@ 1 Introduction
In recent years, large language models have demonstrated remarkable performance on tasks that require reasoning (Bubeck et al., 2023; Wei et al., 2022; Kojima et al., 2022). This has sparked a vigorous debate about the extent to which these models possess reasoning abilities similar to humans (Mitchell & Krakauer, 2023; Mitchell, 2023; Borji, 2023). While proponents argue that reasoning capabilities emerge with scale, referring to LLMs as foundation models (Bommasani et al., 2021; Kaplan et al., 2020), skeptics contend that the models’ performance is a mere reflection of the extensive training data and their vast number of parameters (Razeghi et al., 2022; Wu et al., 2023; Berglund et al., 2024). Thus, the question arises: are these models simply “castles in the air” with “no foundations whatsoever”, as Malik (2021) once stated, or do they possess genuine reasoning capacities? One of the major challenges in this debate is the immense size, complexity, and closed-source nature of popular LLMs and their underlying training data. Moreover, the focus often lies on the models’ performance on downstream reasoning tasks (Fu et al., 2023; Liu et al., 2023), overshadowing in-depth analyses of their reasoning behavior. In this study, we seek to shed light on the ongoing debate by providing a comprehensive overview of research that goes beyond task accuracy, offering more nuanced insights into the models’ reasoning processes. Specifically, we address the following research questions:
@@
While various surveys on reasoning within large language models have emerged in recent years (Huang & Chang, 2023; Yu et al., 2023a; Sun et al., 2024; Yang et al., 2023; Qiao et al., 2023; Chu et al., 2023; Mialon et al., 2023; Liu et al., 2024; Ahn et al., 2024), these studies predominantly focus on techniques that seek to enhance or benchmark the performance of LLMs on downstream reasoning tasks. However, there seems to exist no review of work yet that assesses the reasoning behavior of LLMs, rather than their final task performance. Hence, the contributions of this survey are as follows: (i) To the best of our knowledge, we present the first comprehensive review of literature that evaluates LLM-based reasoning beyond task accuracy, offering deeper insights into the models’ reasoning behavior, and (ii) we suggest a taxonomy that categorizes prevalent evaluation methods designed to analyze the reasoning dynamics of LLMs.
@@ 2 Terminology
We begin this survey with an introduction to fundamental concepts and terminologies pertinent to reasoning in both humans and LLMs. The study of reasoning has captivated scholarly interest for centuries, with insights from diverse disciplines such as philosophy, logic, psychology, neuroscience and artificial intelligence (Holyoak & Morrison, 2005). Over the years, various definitions of reasoning have been proposed, with each discipline offering its unique perspective. In this survey, we define reasoning in the broadest sense (Leighton, 2003; Holyoak & Morrison, 2005; Byrne et al., 1993), and direct the interested reader to the work by Yu et al. (2023a) and Sun et al. (2024) for more domain-specific definitions.
@@ 2.1 A Categorization of Reasoning Tasks
Analogous to the assessment of human reasoning within the field of cognitive psychology (WOOD, 1969; Byrne et al., 1993; Dewey, 2022), the evaluation of reasoning capabilities in large language models predominantly occurs through their engagement with designated reasoning tasks (Sun et al., 2024). These tasks are designed to elicit the system’s capability of drawing conclusions relevant to the problem at hand. In this survey, we distinguish between core and integrated reasoning tasks. Core reasoning tasks are designed to assess fundamental reasoning skills in an isolated manner. They typically involve straightforward applications of a single type of reasoning, such as logical, mathematical or causal reasoning. Examples of such tasks may include syllogisms, basic arithmetic problems or structured causal-graph predictions. Conversely, integrated reasoning tasks require the concurrent use of various reasoning types, thereby assessing a combination of fundamental reasoning skills. Examples are commonsense or scientific reasoning tasks. Such problems often reflect the complex cognitive challenges humans encounter in everyday life and professional settings. A schematic representation distinguishing these task categories is provided in Figure 1.
@@ 3 Reasoning Behavior of LLMs
This section reviews studies that extend beyond mere task accuracy, focusing instead on evaluating the reasoning behavior of large language models. Through this review, we intend to address RQ1 and shed light on how these models currently behave across three core reasoning tasks: logical reasoning (Section 3.1), mathematical reasoning (Section 3.2), and causal reasoning tasks (Section 3.3).
@@ 3.1 Behavior in Logical Reasoning Tasks
The study of logical reasoning evolves around the question of how individuals infer valid conclusions from a set of given premises within a structured framework of logical rules and principles (Holyoak & Morrison, 2005). Based on how conclusions are inferred, logical reasoning can be classified into three types: deductive, inductive, and abductive reasoning (Johnson-Laird, 2008; Flach & Kakas, 2000). While in deductive reasoning, conclusions necessarily follow from the premises’ truth, inductive and abductive reasoning are considered defeasible, i.e. conclusions are at most probable, but never necessary (Koons, 2022). Inductive reasoning is concerned with deriving general conclusions from specific instances, whereas abductive reasoning entails formulating plausible hypotheses that explain the data observed. For additional details on the distinction between the three types of logical reasoning, we refer to prior surveys on reasoning performance within LLMs, such as the ones by Yu et al. (2023a) or Sun et al. (2024).
@@ 3.1.1 Behavior in Deductive Reasoning Tasks
Extensive research has been dedicated to examining the reasoning behavior of LLMs in the context of deductive reasoning tasks. Various studies analyze the validity and consistency of logical proofs generated by LLMs within deductive reasoning problems. For instance, Saparov & He (2023) conduct a systematic assessment of rationales produced by various GPT-3 iterations (Brown et al., 2020), through chain-of-thought (CoT) prompting (Wei et al., 2022). By parsing each step of the generated rationale into first-order-logic, its validity, atomicity, and utility [^1] Validity denotes whether the proof step logically follows from preceding steps, atomicity reflects whether it can be proven with exactly one application of a deduction rule, and utility measures its direct contribution towards the derivation of the final conclusion is measured. Findings indicate that, in comparison to their smaller counterparts, larger models are more adept at generating both valid and atomic steps. However, the utility of these steps is often low, resulting in misleading steps from which the models struggle to recover. In a similar vein, Dziri et al. (2023) study rationales of LLMs by parsing them into computation graphs, where each node represents a partial solution of the multi-step reasoning process. Evaluations on Einstein’s puzzle (Vassberg & Vassberg, 2009) indicate that GPT-3, ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI et al., 2024) solve the task by reducing multi-step reasoning into shortcut pattern matching. This shortcut behavior can yield correct answers when the model has been exposed to similar patterns during training, but falls short in generalizing to novel or more complex instances. Furthermore, theoretical findings suggest that due to the models’ autoregressive nature, mistakes in early stages of the reasoning process can lead to compounding errors that exponentially diminish the likelihood of arriving at the correct solution in subsequent steps.
@@ 3.3 Behavior in Causal Reasoning Tasks
Causal reasoning is the process of discerning the cause-and-effect relationships that govern the dynamics of our environment (Sloman, 2005). Extending beyond correlation, it offers a more nuanced understanding of how changes in one variable can bring about changes in another (Pearl, 2009). While research on the behavior of LLMs in causal reasoning tasks is still in its early stages, the field is receiving growing attention. Various studies to date suggest that LLMs are capable of reciting causal facts from their training data, but lack an intrinsic ability to comprehend or construct causal relationships. For instance, Zečević et al. (2023) indicate that while LLMs, including GPT-3, Meta AI’s Opt, and AlephAlpha’s Luminous (AlephAlpha, 2022), demonstrate some proficiency in causal reasoning tasks that align with causal facts seen during training, they exhibit limited ability to accurately discern and apply causal relationships in scenarios that demand more than associative recalls from their training data. Jin et al. (2023) probe the models’ behavior across three levels of causation: (1) the associational, (2) the interventional, and (3) the counterfactual, as outlined by Pearl & Mackenzie (2018)’s Ladder of Causation. Similarly, their findings indicate that LLMs are more adept at answering associational queries than tackling interventional or counterfactual tasks. Models seem to particularly struggle with causal relationships that deviate from commonsense or are unlikely to be part of their training corpora. Jin et al. (2024) evaluate the ability of LLMs to infer causation from statements that describe correlations between variables. Analyses reveal significant challenges across seventeen LLMs with solving the given task. Although fine-tuning offers substantial improvements, models still fail to generalize in out-of-distribution scenarios. Kosoy et al. (2023)’s evaluation of GPT-3 and PaLM on the blicket detector task, where models need to infer which objects cause a light to switch on (Gopnik & Sobel, 2000), further highlights challenges in causal reasoning with LLMs. While models can identify the correct causal structure when a set of causal hypotheses is provided, they struggle in conditions where the hypotheses are not given, underscoring a limitation in their ability to infer causal relationships from limited data.
@@
Positioned at the last level of Pearl & Mackenzie (2018)’s Ladder of Causation, counterfactual tasks assess the models’ ability to reason about hypothetical scenarios. Research into the behavior of LLMs in counterfactual scenarios reveals notable challenges in current models. Studies, like those conducted by Frohberg & Binder (2022), focus on the ability of LLMs to predict outcomes in hypothetical setups, identifying a significant gap between the capabilities of humans and LLMs, particularly in highly unrealistic scenarios. Li et al. (2023) further illustrate that while models like GPT-3 can produce outcomes that seem to align with counterfactual propositions, these models heavily rely on simple lexical cues within the given context, rather than demonstrating a profound grasp of the scenarios’ hypothetical essence. Yu et al. (2023b) evaluate LLMs on questions embedded in counterfactual presuppositions, pushing the models beyond simple fact retrieval. Their findings suggest that “closed-book” models such as GPT-3 and code-davinci-002 may fabricate facts or base their responses on flawed premises when answering counterfactual questions. Huang et al. (2024) explore a different angle by asking models to modify a given piece of argumentative text so that it upholds a specified logical relationship under new premises. Findings indicate that while LLMs like GPT-3.5 and GPT-4 demonstrate a capacity for solving such tasks, they struggle with modifying arguments such that a new premise makes the argument less likely to be true.
@@
Regarding RQ1, while our review finds that the reasoning behaviors of LLMs are nuanced and diverse, we can discern significant trends: Slight alterations in the task context can markedly impair the models’ reasoning, multi-step reasoning is often reduced to shortcut pattern matching, and scenarios that deviate from the training distribution pose substantial challenges for current models. In addition, fundamental conceptual errors have been identified that highlight the models’ deficiency in understanding basic principles in logic, mathematics and causal reasoning, especially in counterfactual setups.
@@ 4 Evaluation Methods
So far, we have provided an overview of studies that evaluate the behavior of LLMs on three core reasoning tasks. However, to date, a standardized methodology for assessing the reasoning capabilities of large language models is lacking. To address RQ2, we review and categorize predominant evaluation frameworks for analyzing the behavior of LLMs in reasoning tasks. As depicted in Figure 2, we categorize evaluation methodologies into four distinct groups: (i) conclusion-based, (ii) rationale-based, (iii) interactive, (iv) and mechanistic evaluations. In the following, we further delineate each category by discussing prevalent techniques. Additional details can be found in Appendix B.
@@ 4.1 Conclusion-Based Evaluation
In conclusion-based evaluation schemes, the emphasis is placed on the model’s final answer, rather than the process by which the model formulates its conclusion. This outcome-oriented approach, despite its limitation in overlooking the model’s rationales, can nonetheless provide insights into the model’s reasoning behavior. For instance, a thorough error analysis can unveil conceptual errors (Sanyal et al., 2022), cognitive biases (Dasgupta et al., 2022), or a sensitivity to the task context (Wu et al., 2023). Similarly, an examination of the model’s output distribution may reveal inherent predispositions towards certain outcomes (Itzhak et al., 2023), or serve as an indicator of the model’s confidence towards specific conclusions (Frohberg & Binder, 2022). To circumvent the issue of data contamination and test the model’s capacities to generalize, dynamic benchmarks can be employed that update over time. For instance, functional benchmarks capture the underlying reasoning process rather than a static problem formulation, probing the model’s reasoning capabilities on dataset snapshots released every quarter of the year (Srivastava et al., 2024).

### v2

@@ Abstract
Large language models (LLMs) have recently shown impressive performance on tasks involving reasoning, leading to a lively debate on whether these models possess reasoning capabilities similar to humans. However, despite these successes, the depth of LLMs’ reasoning abilities remains uncertain. This uncertainty partly stems from the predominant focus on task performance, measured through shallow accuracy metrics, rather than a thorough investigation of the models’ reasoning behavior. This paper seeks to address this gap by providing a comprehensive review of studies that go beyond task accuracy, offering deeper insights into the models’ reasoning processes. Furthermore, we survey prevalent methodologies to evaluate the reasoning behavior of LLMs, emphasizing current trends and efforts towards more nuanced reasoning analyses. Our review suggests that LLMs tend to rely on surface-level patterns and correlations in their training data, rather than on sophisticated reasoning abilities. Additionally, we identify the need for further research that delineates the key differences between human and LLM-based reasoning. Through this survey, we aim to shed light on the complex reasoning processes within LLMs.
@@ 1 Introduction
In recent years, large language models have demonstrated remarkable performance on tasks that require reasoning (Bubeck et al., 2023; Wei et al., 2022; Kojima et al., 2022). This has sparked a vigorous debate about the extent to which these models possess reasoning abilities similar to humans (Mitchell & Krakauer, 2023; Mitchell, 2023; Borji, 2023). While proponents argue that reasoning capabilities emerge with scale, referring to LLMs as foundation models (Bommasani et al., 2021; Kaplan et al., 2020), skeptics contend that the models’ performance primarily reflects their capacity to memorize the vast amount of data they have been trained on (Wu et al., 2024; Dziri et al., 2023; Razeghi et al., 2022; Zhang et al., 2023). Thus, the question arises: are these models simply “castles in the air” with “no foundations whatsoever,” as Malik (2021) once stated, or do they possess genuine reasoning capacities? One of the major challenges in this debate is the immense size, complexity, and closed-source nature of popular LLMs and their underlying training data. Moreover, the focus often lies on the models’ performance on downstream reasoning tasks (Fu et al., 2023; Liu et al., 2023), overshadowing in-depth analyses of their reasoning behavior. In this study, we seek to shed light on the ongoing debate by providing a comprehensive overview of research that goes beyond task accuracy, offering more nuanced insights into the models’ reasoning processes. Specifically, we address the following research questions:
@@
While various surveys on reasoning within large language models have emerged in recent years (Huang & Chang, 2023; Yu et al., 2024; Sun et al., 2024; Yang et al., 2023; Qiao et al., 2023; Chu et al., 2023; Mialon et al., 2023; Liu et al., 2024; Ahn et al., 2024), these studies predominantly focus on techniques that seek to enhance or benchmark the performance of LLMs on downstream reasoning tasks. However, there seems to exist no review of work yet that assesses the reasoning behavior of LLMs, rather than their final task performance. Hence, the contributions of this survey are as follows: (i) To the best of our knowledge, we present the first comprehensive review of literature that evaluates LLM-based reasoning beyond task accuracy, offering deeper insights into the models’ reasoning behavior, and (ii) we suggest a taxonomy that categorizes prevalent evaluation methods designed to analyze the reasoning dynamics of LLMs.
@@ 2 Terminology
We begin this survey with an introduction to fundamental concepts and terminologies pertinent to reasoning in both humans and LLMs. The study of reasoning has captivated scholarly interest for centuries, with insights from diverse disciplines such as philosophy, logic, psychology, neuroscience and artificial intelligence (Holyoak & Morrison, 2005). Over the years, various definitions of reasoning have been proposed, with each discipline offering its unique perspective. In this survey, we define reasoning in the broadest sense (Leighton, 2003; Holyoak & Morrison, 2005; Byrne et al., 1993), and direct the interested reader to the work by Yu et al. (2024) and Sun et al. (2024) for more domain-specific definitions.
@@ 2.1 A Categorization of Reasoning Tasks
Analogous to the assessment of human reasoning within the field of cognitive psychology (WOOD, 1969; Byrne et al., 1993; Dewey, 2022), the evaluation of reasoning capabilities in large language models predominantly occurs through their engagement with designated reasoning tasks (Sun et al., 2024). These tasks are designed to elicit the system’s capability of drawing conclusions relevant to the problem at hand. In this survey, we distinguish between core and integrated reasoning tasks. Core reasoning tasks are designed to assess fundamental reasoning skills in an isolated manner. They typically aim to test a single type of reasoning, such as logical, mathematical or causal reasoning. Examples of such tasks may include syllogisms, basic arithmetic problems or structured causal-graph predictions. Conversely, integrated reasoning tasks require the concurrent use of various reasoning types, thereby assessing a combination of fundamental reasoning skills. Examples are commonsense or scientific reasoning tasks. Such problems often reflect the complex cognitive challenges humans encounter in everyday life and professional settings. A schematic representation distinguishing these task categories is provided in Figure 1.
@@ 3 Reasoning Behavior of LLMs
This section reviews studies that extend beyond mere task accuracy, focusing instead on evaluating the reasoning behavior of large language models. [^1] We consistently specify when a particular technique, other than standard prompting, is utilized. For a review of prompting approaches designed to improve the models’ reasoning performance, we recommend consulting prior surveys by Huang & Chang (2023),  Yu et al. (2024), or  Sun et al. (2024). Through this review, we intend to address RQ1 and shed light on how these models currently behave across three core reasoning tasks: logical reasoning (Section 3.1), mathematical reasoning (Section 3.2), and causal reasoning (Section 3.3).
@@ 3.1 Behavior in Logical Reasoning Tasks
The study of logical reasoning evolves around the question of how individuals infer valid conclusions from a set of given premises within a structured framework of logical rules and principles (Holyoak & Morrison, 2005). Based on how conclusions are inferred, logical reasoning can be classified into three types: deductive, inductive, and abductive reasoning (Johnson-Laird, 2008; Flach & Kakas, 2000). While in deductive reasoning, conclusions necessarily follow from the premises’ truth, inductive and abductive reasoning are considered defeasible, i.e. conclusions are at most probable, but never necessary (Koons, 2022). Inductive reasoning is concerned with deriving general conclusions from specific instances, whereas abductive reasoning entails formulating plausible hypotheses that explain the data observed. For additional details on the distinction between the three types of logical reasoning, we refer to prior surveys on reasoning performance within LLMs, such as the ones by Yu et al. (2024) or Sun et al. (2024).
@@ 3.1.1 Behavior in Deductive Reasoning Tasks
Extensive research has been dedicated to examining the reasoning behavior of LLMs in the context of deductive reasoning tasks. Various studies analyze the validity and consistency of logical proofs generated by LLMs within deductive reasoning problems. For instance, Saparov & He (2023) conduct a systematic assessment of the rationales produced by various GPT-3 iterations (Brown et al., 2020), through chain-of-thought (CoT) prompting (Wei et al., 2022). By parsing each step of the generated reasoning trace into first-order-logic, its validity, atomicity, and utility [^2] Validity denotes whether the proof step logically follows from preceding steps, atomicity reflects whether it can be proven with exactly one application of a deduction rule, and utility measures its direct contribution towards the derivation of the final conclusion. is measured. Findings indicate that, in comparison to their smaller counterparts, larger models are more adept at generating both valid and atomic steps. However, the utility of these steps is often low, resulting in misleading steps from which the models struggle to recover. In a similar vein, Dziri et al. (2023) study rationales of LLMs by parsing them into computation graphs, where each node represents a partial solution of the multi-step reasoning process. Evaluations on Einstein’s puzzle (Vassberg & Vassberg, 2009) indicate that GPT-3, ChatGPT (OpenAI, 2022) and GPT-4 (OpenAI et al., 2024) solve the task by reducing multi-step reasoning into shortcut pattern matching. This shortcut behavior can yield correct answers when the model has been exposed to similar patterns during training, but falls short in generalizing to novel or more complex instances. Furthermore, theoretical findings suggest that due to the models’ autoregressive nature, mistakes in early stages of the reasoning process can lead to compounding errors that exponentially diminish the likelihood of arriving at the correct solution in subsequent steps.
@@ 3.3 Behavior in Causal Reasoning Tasks
Causal reasoning is the process of discerning the cause-and-effect relationships that govern the dynamics of our environment (Sloman, 2005). Extending beyond correlation, it offers a more nuanced understanding of how changes in one variable can bring about changes in another (Pearl, 2009). For a comprehensive introduction to causal reasoning, we strongly recommend the work of Pearl (2009). While research on the behavior of LLMs in causal reasoning tasks is still in its early stages, the field is receiving growing attention. Various studies to date suggest that LLMs are capable of reciting causal facts from their training data, but lack an intrinsic ability to comprehend or construct causal relationships. For instance, Zečević et al. (2023) indicate that while LLMs, including GPT-3, Meta AI’s Opt, and AlephAlpha’s Luminous (AlephAlpha, 2022), demonstrate some proficiency in causal reasoning tasks that align with causal facts seen during training, they exhibit limited ability to accurately discern and apply causal relationships in scenarios that demand more than associative recalls from their training data. Jin et al. (2023) probe the models’ behavior across three levels of causation: (1) the associational, (2) the interventional, and (3) the counterfactual, as outlined by Pearl & Mackenzie (2018)’s Ladder of Causation. Similarly, their findings indicate that LLMs are more adept at answering associational queries than tackling interventional or counterfactual tasks. Models seem to particularly struggle with causal relationships that deviate from commonsense or are unlikely to be part of their training corpora. Jin et al. (2024) evaluate the ability of LLMs to infer causation from statements that describe correlations between variables. Analyses reveal significant challenges in solving the given task across seventeen LLMs. Although fine-tuning offers substantial improvements, models still fail to generalize in out-of-distribution scenarios. Kosoy et al. (2023)’s evaluation of GPT-3 and PaLM on the blicket detector task (Gopnik & Sobel, 2000), where models need to infer which objects cause a light to switch on, further highlights challenges in causal reasoning with LLMs. While models can identify the correct causal structure when a set of causal hypotheses is provided, they struggle in conditions where the hypotheses are not given, underscoring a limitation in their ability to infer causal relationships from limited data.
@@
Positioned at the last level of Pearl & Mackenzie (2018)’s Ladder of Causation, counterfactual tasks assess the models’ ability to reason about hypothetical scenarios. Research into the behavior of LLMs in counterfactual scenarios reveals notable challenges in current models. Studies, like those conducted by Frohberg & Binder (2022), focus on the ability of LLMs to predict outcomes in hypothetical setups, identifying a significant gap between the capabilities of humans and LLMs, particularly in highly unrealistic scenarios. Li et al. (2023) further illustrate that while models like GPT-3 can produce outcomes that seem to align with counterfactual propositions, these models heavily rely on simple lexical cues within the given context, rather than demonstrating a profound grasp of the scenarios’ hypothetical essence. Yu et al. (2023) evaluate LLMs on questions embedded in counterfactual presuppositions, pushing the models beyond simple fact retrieval. Their findings suggest that “closed-book” models such as GPT-3 and code-davinci-002 may fabricate facts or base their responses on flawed premises when answering counterfactual questions. Huang et al. (2024) explore a different angle by asking models to modify a given piece of argumentative text so that it upholds a specified logical relationship under new premises. Findings indicate that while LLMs like GPT-3.5 and GPT-4 demonstrate a capacity for solving such tasks, they struggle with modifying arguments such that a new premise makes the argument less likely to be true, an observation also made by Rudinger et al. (2020) in the context of abductive reasoning.
@@
Regarding RQ1, our review suggests that the reasoning behaviors of LLMs are nuanced and diverse, yet a significant trend emerges: while current LLMs demonstrate proficiency in reasoning problems that align with their training distribution, they frequently encounter substantial conceptual difficulties in out-of-distribution scenarios. Notably, slight alterations in the task context can markedly impair the models’ reasoning capabilities. Multi-step reasoning is often reduced to shortcut pattern matching, and fundamental conceptual errors highlight the models’ deficiency in understanding basic principles of logic, mathematics, and causal reasoning, particularly in counterfactual setups.
@@ 4 Evaluation Methods
So far, we have provided an overview of studies that evaluate the behavior of LLMs on three core reasoning tasks. However, to date, a standardized methodology for assessing the reasoning capabilities of large language models is lacking. To address RQ2, we review and categorize predominant evaluation frameworks for analyzing the behavior of LLMs in reasoning tasks. As depicted in Figure 2, we categorize evaluation methodologies into four distinct groups: (i) conclusion-based, (ii) rationale-based, (iii) interactive, (iv) and mechanistic evaluations. In the following sections, we further delineate each category by discussing prevalent techniques. An overview of the advantages and disadvantages of each evaluation procedure can be found in Table 1. Additional details are provided in Appendix B.
@@ 4.1 Conclusion-Based Evaluation
In conclusion-based evaluation schemes, emphasis is placed on the model’s final answer rather than the process by which the model arrives at its conclusion. This outcome-oriented approach, despite its limitation of overlooking the model’s rationales, can nonetheless provide insights into the model’s reasoning behavior. For instance, a thorough error analysis can unveil conceptual errors (Sanyal et al., 2022), reflections of cognitive biases (Dasgupta et al., 2022), or sensitivities to the task context (Wu et al., 2024). Similarly, an examination of the model’s output distribution may reveal inherent predispositions towards certain outcomes (Itzhak et al., 2024), or serve as an indicator of the model’s confidence in specific conclusions (Frohberg & Binder, 2022). However, relying solely on the model’s final conclusion can be less reliable than also considering how the model arrives at its conclusion. For example, candidate answers derived from first token probabilities in multiple-choice setups are often not robust (Wang et al., 2024b), and final answers might not always align with the model’s verbalized reasoning (Mondorf & Plank, 2024; Ye & Durrett, 2022). Moreover, answers to benchmark datasets might have been compromised by data leakage during the model’s training process, limiting the insights that can be drawn from a correct conclusion (Balloccu et al., 2024; Xu et al., 2024). To address the issue of data contamination and test the model’s capacity to generalize, dynamic benchmarks can be employed that update over time. For instance, functional benchmarks capture the underlying reasoning process rather than a static problem formulation, probing the model’s reasoning capabilities on dataset snapshots released quarterly (Srivastava et al., 2024).